40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 14063     Entry time: Fri Jul 13 02:52:11 2018     In reply to: 14061     Reply to this: 14067   14068
Author: gautam 
Type: Update 
Category: General 
Subject: Vent objectives and prep 

[KA, GV]

  1. Arm scan setup moved from West side of PSL table to North side of AP table
    • Marconi is now providing the LO for the loop. A ~5m 50ohm BNC cable needs to be laid out from the PLL area to the NW corner of the AP table where the Agilent is for the TF measurement scheme which allows phase discrimination.
    • We took this opportunity to characterize/improve the setup a bit, and also clean up the 10s of cables around the PSL table. There were some Video cables (75ohm instead of 50ohm no) that were part of the scanning setup which we excised.
    • Mode matching onto beat PD on PSL table was improved - beat signal Vpp increased by factor of 2. PLL gain was adjusted accordingly. 
    • AUX power @AP table ~37mW (before recombination BS), ~3.7mW onto SRM, ~200uW onto ITMY.
    • We got a beatnote in transmission of ~120uV (rms?) after optimizing alignment into Y arm cavity with the AUX frequency on an arm FSR.
    • From Sandrine's calculations, I expected ~20mV of beat signal. We leave it to Annalisa and Terra to square that circle. Probably the MM into the arm isn't stellar either, and we didn't check the polarization of the aux beam (that it is matched to the IFO's p-pol).
    • The entire AUX injection chain needs careful characterization before
  2. Vent prep
    • Both arms were locked to IR, TRY maximized using ASS, TRX maximized by hand.
    • GTRY and GTRX were also maximized.
    • All test mass/PRM/SRM oplevs were centered with this "good" alignment.
    • PSL power into the IMC was cut from 1.07 W (measured after G&H mirror) to 97 mW, by rotating the waveplate immediately after the PSL (original angle 180deg, new angle 216 degrees).
    • PMC was locked. I did not need to change any of the gain settings.
    • 2" R=10% BS in the IMC REFL path was replaced with a Y2, so there is no MCREFL till we turn the power back up.
    • IMC was locked. I touched up my IMC low power autolocker, but this probably needs a bit more touching up tomorrow. MC has remained locked at low power for ~25mins, but as I typed this, I jinxed it and it lost the lock. anyway we have good alignment references.
    • PSL shutter will remain closed.

@SV, we are ready to vent tomorrow. Aaron is supposed to show up ~830am to assist.

Attachment 1: preVentStatus.png  220 kB  Uploaded Fri Jul 13 04:10:30 2018  | Hide | Hide all
preVentStatus.png
ELOG V3.1.3-